# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ C.1.3 Line 2 to 2: Conditions on Phase Transition
There are three possible conditions to end the exploration period in a phase $k$. The first one is Line 2, judge whether the algorithm should enter the exploitation period. Intuitively speaking, the inequality $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ provides a recommended qualified arm with confidence $1-\Theta(\delta_{k})$. Since $A_{t}^{\text{ee}}$ is the action we take in the last round, this condition also guarantees $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds before the collecting the latest sample of $A_{t}^{\text{ee}}$. To further confirm whether arm $A_{t}^{\text{ee}}$ is qualified, we need to keep pulling arm $A_{t}^{\text{ee}}$ by collecting samples independent of the exploration periods. Line 2 suggests that the algorithm will terminate and output $A_{t}^{\text{ee}}$, if the LCB calculated by the tolerance level $\delta/\alpha_{k}$ is above $\mu_{0}$ during the exploitation process. Before entering the while loop (Line 2) in the exploitation period, Line 2 transfer the latest collected sample of $A_{t}^{\text{ee}}$ in the exploration period into $Q$. Usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Similar to the reason described in the section 4.1 of Li and Cheung (2025), Lemma 22 is required to prove Lemma 23, further Theorem 3. We put more discussion of $Q$ to section C.1.4.
@@
The third condition is Line 2, which is correlated to the budget of the exploration period. For a phase $k$, Algorithm 2 sets up a maximum pulling times $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ for each arm $a\in B$. Recall we only consider picking $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2. The third condition is Line 2 is to enter the exploration period in phase $k+1$, if pulling times are no less than $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1$ holds for all $a\in B$. Line 2 and 2 also guarantee Algorithm 2 will not get stuck into a infinite loop of “While True” in Line 2.
@@ C.1.4 Line 2 to 2: Exploration
As we mentioned in section C.1.3, usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Someone may question whether we can directly abandon the latest collected sample of $A_{t}^{\text{ee}}$ in $\mathcal{H}$, which seems also able to guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. We cannot to do so. By transferring samples into or out $Q$, we are able to derive
@@
The reason is we do not change the arrival order of $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$ for each $a\in B$. To analyze Algorithm 2, we need to introduce
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. More details are in section C.2. Lemma 13 suggests that for phase $k\geq\kappa^{\text{ee}}_{b}$, we can guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{a}<\text{UCB}_{a}^{\text{ee}}$ holds for all $a\in B_{b}$, which further helps to prove upper bounds of the pulling times for all $a\in B_{b}$. Simply dropping the latest collected samples are unable to guarantee 13, making it much harder to conduct the analysis.
@@
The reason is all the collected samples in the exploration periods are either in $\mathcal{H}^{\text{ee}}$ or $Q$. Notice that we only transfer sample from $\mathcal{H}^{\text{ee}}$ to $Q$ in Line 2. In this case, $A_{t}^{\text{ee}}$ just get pulled in the last round of exploration. Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in 2 guarantees Algorithm 2 will never collect new samples of arm $a$ in the exploration period. This property implies following.
@@
At the end of this subsection, we present a Lemma similar to 13, but for the exploitation period.
@@ C.1.5 Adaptation of History Literature
Compared to the Algorithm SEE in Li and Cheung (2025), our main adaptation consists of two points. The first adaptation is to put the stopping conditions (Line 2 to 2) before the exploration periods (Line 2 to 2). This design is to guarantee Lemma 29, guaranteeing Algorithm 2 will not pull arm $a$ whose $\text{UCB}_{a}^{\text{ee}}$ is below $\mu_{0}$.
The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, instead of maximum pulling times of the total pulling times (reference line index is to be added). The reason is the arm in a bracket $B$ is randomized, it is hard to analyze the required pulling times that a bracket need to output a $\hat{a}_{k}$.
@@ C.2 Properties of Algorithm 2
This subsections follows the Appendix B in Li and Cheung (2025), which are nearly the same. Here we remain the full proof for completeness, as we have modified some implementation details and proof.
@@
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. The above definition of $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are conditioned on the realized reward of $\{X_{a,s}^{\text{ee}}\}_{a=1,s=1}^{K,+\infty}$, $\{X_{a,s}^{\text{et}}\}_{a=1,s=1}^{K,+\infty}$ collected by the algorithm copy $b$. $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are the minimum phase index such that concentration event holds. Since $B_{b}$ is determined by an exogenous random permutation, we know $B_{b}$ is independent with $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$.
@@
The first step is still caused by the Line 2 in Algorithm 2. The second step is from the foundation. The last line shows that the induction holds. ∎
@@
$=$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
@@
We apply induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
If the conclusion holds for the phase k1k-1, we prove the case of $k$ by discussing all the possible values of $\hat{a}_{k-1}$. Before the start of phase $k$, there are three possible scenarios for the value of $\hat{a}_{k-1}$. Phase k1k-1 may end with taking $\hat{a}_{k-1}$ takes value at the Line 2, 2 or 2.
@@
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, We can also assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$. Prove by contradiction. Assume there exists $a\in[K]$ such that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})\geq\mu_{0}$ at the end of phase k1k-1. From the induction, since $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ at the start of phase k1k-1, the algorithm would enter Line 2 and $\hat{a}_{k-1}$ cannot take value Not Completed. Thus, we can assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$ holds at the end of phase k1k-1. Further, at the start of phase $k$, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$.
The remaining work is to prove the conclusion in the case that $\hat{a}_{k-1}\in[K]$ at the Line 2. We can first assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\neq\hat{a}_{k-1}$. The reason is from the induction, $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ holds for all $a$ at the start of phase k1k-1. Once there exists an arm $a$ whose LCB is above $\mu_{0}$ after an arm pull, the algorithm will enter Line 2, suggesting that there is at most one arm whose LCB is above $\mu_{0}$.
@@
Since we put the last collected sample $X$ into $Q$ in Line 2, we can assert $\hat{\mu}_{\hat{a}_{k-1}}$ before the start of phase $k$ must be $\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}$. Since $U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k}}{K}\right)>U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)$, we know before the start of phase $k$, $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ must hold.
@@
The remaining work is to prove if $\hat{a}_{k}\in[K]$, we have $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$. For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$, and prove by contradiction. If $\hat{a}\in[K]$ and $\mu_{\hat{a}}<\omega\mu_{j}+(1-\omega)\mu_{0}<\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}$, we have
$A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})$
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.$
@@
As we take $U(t,\frac{\delta}{K})=\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2K(\lceil\log_{2}t\rceil)^{2}}{\delta}}}{t}$, we get
@@
By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, which means
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
@@
we can conclude
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$
@@
$\Leftarrow$ $\omega\mu_{1}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
@@
$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$
In summary, we can conclude
@@
The last line suggests a contradiction. We can conclude $\mu_{\hat{a}}\geq\frac{C-1}{C+3}\mu_{j}+\frac{4}{C+3}\mu_{0}$. ∎
@@
hols for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If the above conclusion holds, we can use induction to conclude
@@
This is a direct conclusion from Lemma 25, 26. ∎

### v2

@@ C.1.3 Line 2 to 2: Conditions on Phase Transition
There are three possible conditions to end the exploration period in a phase $k$. The first one is Line 2, judge whether the algorithm should enter the exploitation period. Intuitively speaking, the inequality $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ provides a recommended qualified arm with confidence $1-\Theta(\delta_{k})$. Since $A_{t}^{\text{ee}}$ is the action we take in the last round, this condition also guarantees $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds before the collecting the latest sample of $A_{t}^{\text{ee}}$. To further confirm whether arm $A_{t}^{\text{ee}}$ is qualified, we need to keep pulling arm $A_{t}^{\text{ee}}$ by collecting samples independent of the exploration periods. Line 2 suggests that the algorithm will terminate and output $A_{t}^{\text{ee}}$, if the LCB calculated by the tolerance level $\delta/\alpha_{k}$ is above $\mu_{0}$ during the exploitation process. Before entering the while loop (Line 2) in the exploitation period, Line 2 transfer the latest collected sample of $A_{t}^{\text{ee}}$ in the exploration period into $Q$. Usage of $Q$ is to guarantee Lemma 23, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Similar to the reason described in the section 4.1 of Li and Cheung [2025], Lemma 23 is required to prove Lemma 24, further Theorem 3. We put more discussion of $Q$ to section C.1.4.
@@
The third condition is Line 2, which is correlated to the budget of the exploration period. In phase $k$, each arm $a\in B$ is assigned an exploration budget of $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$. We only consider pull $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2. If all arms in $B$ have reached their budget (up to the 1-1 offset in the implementation), the algorithm sets $\hat{a}_{k}=\textsf{Not Completed}$ and transitions to phase $k+1$. Together, Lines 2 and 2 ensure that the while-loop in Line 2 cannot become an infinite loop.
@@ C.1.4 Line 2 to 2: Exploration
As noted above, the purpose of $Q$ is to enforce Lemma 23, which requires $\mathrm{LCB}^{\mathrm{ee}}_{a}<\mu_{0}$ for all $a\in B$ at the start of each phase. One might ask whether it suffices to simply discard the most recent exploration sample instead of storing it in $Q$. This is not sufficient. The transfer mechanism preserves the per-arm arrival order of samples, which is needed for the following property.
@@
The key point is that the algorithm never permutes the order of $\{X_{a,s}^{\mathrm{ee}}\}_{s\geq 1}$ for any arm $a\in B$; it only temporarily holds out the most recent sample and later reinserts it without changing order.
For the analysis, we introduce the random phase indices
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}.$
$\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. More details are in section C.2. Lemma 14 suggests that for phase $k\geq\kappa^{\text{ee}}_{b}$, we can guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{a}<\text{UCB}_{a}^{\text{ee}}$ holds for all $a\in B_{b}$, which further helps to prove upper bounds of the pulling times for all $a\in B_{b}$. Simply dropping the latest collected samples is unable to guarantee 14, making it much harder to conduct the analysis.
@@
The reason is all the collected samples in the exploration periods are either in $\mathcal{H}^{\text{ee}}$ or $Q$. Notice that we only transfer sample from $\mathcal{H}^{\text{ee}}$ to $Q$ in Line 2. In this case, $A_{t}^{\text{ee}}$ just get pulled in the last round of exploration. Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in Line 2 ensures that Algorithm 2 will not draw a fresh exploration sample for arm $a$ until that held-out sample is transferred back. This yields the following bound.
@@
At the end of this subsection, Lemma 17 is similar to 14, but for the exploitation period.
@@ C.1.5 Adaptation of History Literature
Compared to the Algorithm SEE in Li and Cheung [2025], our main adaptation consists of two points. The first adaptation is we place the stopping/transition conditions (Lines 22) before the exploration step (Lines 22). This ordering is required for Lemma 30, which ensures that the algorithm does not sample any arm $a$ whose exploration upper confidence bound satisfies $\mathrm{UCB}^{\mathrm{ee}}_{a}\leq\mu_{0}$.
The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, rather than a budget on the total number of exploration pulls across the bracket (Line 2 and 2). This modification is motivated by the randomization of arms within each bracket: with randomized arm assignments, it becomes difficult to track how a total budget is required for the bracket to produce a candidate $\hat{a}_{k}$.
@@ C.2 Properties of Algorithm 2
This subsection closely follows Appendix B of Li and Cheung [2025], with only minor differences. Since we modify several implementation details (and correspondingly adjust parts of the argument), we include the full proofs for completeness.
@@
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}.$
$\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. The above definition of $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are conditioned on the realized reward of $\{X_{a,s}^{\text{ee}}\}_{a=1,s=1}^{K,+\infty}$, $\{X_{a,s}^{\text{et}}\}_{a=1,s=1}^{K,+\infty}$ collected by the algorithm copy $b$. $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are the minimum phase index such that concentration event holds. Since $B_{b}$ is determined by an exogenous random permutation, we know $B_{b}$ is independent with $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$.
@@
The first step is still caused by the Line 2 in Algorithm 2. The second step is from the conclusion at phase index $k$. The last line shows that the induction holds. ∎
@@
$=$ $\sum_{a=1}^{K}\left(N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\right)\mathds{1}(\hat{a}_{k}=a)$
@@
We prove the Lemma by induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
If the conclusion holds for the phase k1k-1, we prove the case of $k$ by discussing all the possible values of $\hat{a}_{k-1}$. Since the pulling process enter phase $k$, we can conclude phase k1k-1 may end with taking $\hat{a}_{k-1}$ at the Line 2, 2 or 2.
@@
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, We can also assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$. We argue by contradiction. Assume there exists $a\in[K]$ such that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})\geq\mu_{0}$ at the end of phase k1k-1. From the induction, since $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ at the start of phase k1k-1, the algorithm would enter Line 2 and $\hat{a}_{k-1}$ cannot take value Not Completed. Thus, we can assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$ holds at the end of phase k1k-1. Further, at the start of phase $k$, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$.
The remaining work is to prove the conclusion in the case that $\hat{a}_{k-1}\in[K]$ at the Line 2. We can first assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\neq\hat{a}_{k-1}$. The reason is from the induction, $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ holds for all $a$ at the start of phase k1k-1. Once there exists an arm $a$ whose LCB is above $\mu_{0}$ after an arm pull, the algorithm will enter Line 2, suggesting that there is at most one arm whose LCB is above $\mu_{0}$. This observation guarantees that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$.
@@
Since we put the last collected sample $X$ into $Q$ in Line 2, we can assert $\hat{\mu}_{\hat{a}_{k-1}}$ before the start of phase $k$ must be $\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}$. Since $U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k}}{K}\right)>U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)$, we know before the start of phase $k$, $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ must hold. ∎
@@
The remaining work is to prove if $\hat{a}_{k}\in[K]$, we have $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$. For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$. From the outputting rule, we have
$\begin{split}&A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\\ &\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}\\ &\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.\end{split}$ (25)
@@
We consider two cases, $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})=1$ or $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})\geq 2$. If $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})=1$, we have
$\mu_{0}\leq\hat{\mu}_{\hat{a}_{k},1}-C\cdot U(1,\frac{\delta_{k}}{K})$
$\Rightarrow$ $\mu_{0}\leq\mu_{\hat{a}_{k}}-(C-1)\cdot U(1,\frac{\delta_{k}}{K})$
$\Rightarrow$ $\mu_{0}\leq\mu_{\hat{a}_{k}}-2(C-1)$
$\Leftrightarrow$ $\mu_{0}+2(C-1)\leq\mu_{\hat{a}_{k}}.$
The second last line is from the definition of (22), suggesting that $U(1,\frac{\delta_{k}}{K})\geq 2$. From the assumption $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ and $C+1$, we know $\omega\mu_{j}+(1-\omega)\mu_{0}=\frac{C-1}{C+3}\mu_{j}+(1-\frac{C-1}{C+3})\mu_{0}<\frac{C-1}{4}+\mu_{0}<\mu_{0}+2(C-1)$. We complete the proof.
If $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})\geq 2$, we argue by contradiction. Assume $\mu_{\hat{a}_{k}}<\omega\mu_{j}+(1-\omega)\mu_{0}$. As we take $U(t,\frac{\delta}{K})=\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2K(\lceil\log_{2}t\rceil)^{2}}{\delta}}}{t}$, we get
@@
By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, we have
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$ (26)
@@
Together with (26), we can conclude
$\Rightarrow$ $\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$ (27)
@@
$\Leftarrow$ $\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
@@
$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$ (28)
Combining (27) and (28), we can conclude
@@
[…]
@@
holds for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If (29) holds, we can use induction to conclude
@@
This is a direct conclusion from Lemma 26, 27. ∎